OpenStudy (annipuppi):

if a/4=b/7 what is the value of a/b? Show work

OpenStudy (annipuppi):

@phi

OpenStudy (raden):

a/4 = b/7 first multiply by 1/b on both sides, get a/4 * 1/b = b/7 * 1/b 1/4 * a/b = 1/7 then multiply both sides by 4, what do you get ......

OpenStudy (annipuppi):

1*4a/b=4/7

OpenStudy (annipuppi):

@RadEn

OpenStudy (phi):

\[ \frac{a}{4} = \frac{b}{7} \] if you multiply by 1/b like RadEn did you do this \[ \frac{1}{b} \cdot \frac{a}{4} = \frac{1}{b} \cdot\frac{b}{7} \] on the right side of the =, you have a "b" "up top" and a "b" down below in the fraction that means you have b/b (b divided by itself). You get 1. in other words you get \[ \frac{1}{b} \cdot \frac{a}{4} = \frac{1}{7} \] now multiply both sides by 4: \[ 4 \cdot \frac{1}{b} \cdot \frac{a}{4} = 4 \cdot \frac{1}{7} \]

OpenStudy (phi):

notice on the left side you have a 4 up top and a 4 down below (4 divided by 4) what does 4/4 turn into ?

OpenStudy (annipuppi):

I got the answer. its a/b=4/7 right?

OpenStudy (phi):

yes

OpenStudy (annipuppi):

Okay thanks

OpenStudy (phi):

yw